OpenStudy (anonymous):

what values for theta(0 less than or equal to theta less than or equal to 2 pi) satisfy the equation? 3sin theta = sin theta -1

OpenStudy (anonymous):

This? \(0\leq \theta\leq2\pi\)\[ 3\sin\theta = \sin(\theta-1) \]

Miracrown (miracrown):